### v1

@@ Abstract
Ground-based remote sensing cloud image sequence extrapolation is a key research area in the development of photovoltaic power systems. However, existing approaches exhibit several limitations: (1) they primarily rely on static kernels to augment feature information, lacking adaptive mechanisms to extract features at varying resolutions dynamically; (2) temporal guidance is insufficient, leading to suboptimal modeling of long-range spatiotemporal dependencies; and (3) the quadratic computational cost of attention mechanisms is often overlooked, limiting efficiency in practical deployment. To address these challenges, we propose USF-Net, a Unified Spatiotemporal Fusion Network that integrates adaptive large-kernel convolutions and a low-complexity attention mechanism, combining temporal flow information within an encoder-decoder framework. Specifically, the encoder employs three basic layers to extract features. Followed by the unified spatiotemporal module, which comprises: (1) a spatial information branch equipped with a spatial selection module that dynamically captures multi-scale contextual information, and (2) a temporal information branch featuring a temporal agent attention module that effectively models long-range temporal dependencies while maintaining computational efficiency. In addition, a dynamic spatiotemporal module with a temporal guidance module is introduced to enable unified modeling of temporally guided spatiotemporal dependencies. On the decoder side, a dynamic update module is employed to address the common “ghosting effect.” It utilizes the initial temporal state as an attention operator to preserve critical motion signatures. As a key contribution, we also introduce and release the ASI-Cloud Image Sequence (ASI-CIS) dataset, a new large-scale, high-resolution benchmark designed to address the critical limitations of existing public datasets. Extensive experiments on ASI-CIS demonstrate that USF-Net significantly outperforms state-of-the-art methods, establishing a superior balance between prediction accuracy and computational efficiency for ground-based cloud extrapolation. The dataset and source code will be available at https://github.com/she1110/ASI-CIS.
@@
[…]
@@ 1 Introduction
With the global energy landscape undergoing an accelerated transition toward cleaner sources, photovoltaic (PV) power generation has emerged as one of the fastest-growing sectors in renewable energy due to its advantages of zero-carbon emissions and widespread resource availability [Intro1]. However, the inherent intermittency of solar power, driven by its high dependency on solar irradiance, presents a significant challenge to grid stability. Rapid fluctuations in power output, often caused by fast-moving clouds, can impose substantial pressure on power dispatch systems and energy storage configurations, complicating grid integration [Intro2]. Consequently, the ability to accurately forecast solar irradiance at high temporal resolutions has become a critical enabler for enhancing PV integration capacity, ensuring power system reliability, and facilitating the large-scale deployment of solar technologies.
@@
The output power of PV systems is strongly correlated with solar irradiance, which is primarily modulated by atmospheric factors like cloud cover [Intro3]. Irradiance fluctuations due to cloud obstruction can be broadly categorized into two types: sustained shading (e.g., stratiform cloud coverage) and transient shading (e.g., rapid cumulonimbus movement). While numerical weather prediction (NWP) models can forecast sustained events, transient shading requires real-time, minute-level monitoring to capture its impact [Intro4]. Cloud observation is typically performed using satellite or ground-based remote sensing [Intro5, Intro6]. Satellite-based cloud imagery, with its low spatial resolution (typically >> 1 km) and infrequent updates (typically $\geq$ 30-minute intervals), is inadequate for tracking the localized, dynamic evolution of clouds that cause rapid irradiance changes. In contrast, ground-based cloud imagers provide high-resolution, high-frequency data on cloud structure and dynamics, making them the optimal data source for this task [25TIM]. Therefore, improving the prediction accuracy of ground-based cloud image sequences is critical for achieving precise, ultra-short-term irradiance forecasting, which plays a pivotal role in ensuring the operational stability of grid-connected PV systems [Intro7, Intro8, Intro9].
@@
The task of ground-based remote sensing cloud image sequence extrapolation is fundamentally a deterministic spatiotemporal prediction problem. Advancing the state-of-the-art (SOTA) requires addressing two core technical challenges: a) developing models with the capacity to represent the complex, nonlinear, and multi-scale morphological variations of clouds, and b) achieving this with algorithms that can efficiently capture long-range spatiotemporal dependencies to meet the real-time inference demands of practical applications.
@@
Existing methodologies have sought to address these challenges through two primary avenues: traditional frameworks and, more recently, deep learning-based approaches. Traditional frameworks, which often rely on linear models embedded with vector field representations (e.g., optical flow [Intro10], similarity maximization [Intro11]). However, these methods often suffer from excessive computational overhead and insufficient nonlinear feature representation capabilities, resulting in substantial performance degradation under complex atmospheric conditions.
In response, deep learning techniques have garnered significant attention in spatiotemporal sequence prediction research. Numerous studies have highlighted the effectiveness of recurrent neural networks (RNNs) [Intro12], a classical deep learning algorithm for sequential data processing, in modeling spatiotemporal features. Subsequently, long short-term memory (LSTM) networks [Intro14], as RNN variants, have been successfully applied to capture the temporal dynamics of cloud motion due to their ability to maintain long-range temporal dependencies [Intro13, PredRNN, PredRNN++]. To mitigate the inherent sequential computation constraints of LSTM-based methods, which limit parallelization efficiency, recent works have proposed hybrid architectures integrating convolutional neural networks (CNNs) with LSTM to enhance local spatial feature extraction [ConvLSTM]. Addressing the multi-scale variability inherent in ground-based cloud images, researchers have further developed multi-scale convolutional kernels to extract cloud morphological features across spatial resolutions hierarchically [Intro16, Intro17]. Some innovations like the Motion-Aware Unit (MAU) and various attention mechanisms have been introduced to capture motion patterns and long-range dependencies [MAU, Intro18, Intro19].
While these methods have yielded significant performance gains, critical limitations persist. Firstly, the reliance on static, fixed-size kernels fails to dynamically adapt receptive fields to the diverse and evolving scales of cloud structures, limiting the ability to resolve scale ambiguities in rapidly changing systems. Moreover, the interaction between temporal and spatial feature streams remains underdeveloped, lacking an explicit mechanism for temporal information to guide spatial feature learning. This decoupled approach compromises the capture of long-range relationships. Furthermore, the quadratic complexity of standard self-attention of these solutions is frequently overlooked, becoming a prohibitive bottleneck for the high-resolution imagery and real-time inference required in practical cloud monitoring applications.
The extrapolation of ground-based cloud image sequences is a critical task for enhancing ultra-short-term PV power forecasting. However, existing datasets used in current research suffer from low spatial resolution and hardware-induced visual obstructions, which hinder accurate predictions of PV power generation [MSTANet, CloudPredRNN++]. This limitation primarily stems from the fact that low resolution and hardware obstructions introduce artifact interference, significantly restricting the application of cloud image sequence extrapolation in accurate, real-world PV power forecasting. Consequently, it is particularly important to develop benchmark datasets and techniques for high-resolution, multi-scale, and cross-seasonal cloud image sequence extrapolation.
@@
To address these distinct challenges, this article introduces the Unified Spatiotemporal Fusion Network (USF-Net), a novel architecture that integrates adaptive large-kernel convolutions with a low-complexity attention mechanism within a unified encoder-decoder framework. Specifically, the encoder employs depthwise separable (DW) convolutions and squeeze-excitation (SE) blocks for hierarchical downsampling and multi-scale feature extraction. Following the encoder, we propose a unified spatiotemporal module (USTM) comprising three core components: 1) a spatial information branch, in which a spatial selection module (SSM) is designed to extract multi-scale cloud context through adaptive receptive field adjustment dynamically; 2) a temporal information branch incorporating a temporal agent attention module (TAM) is introduced to capture long-range temporal dependencies in cloud sequences while reducing computational complexity; and 3) a dynamic spatiotemporal module: a temporal guidance module (TGM) fuses spatial and temporal features, enabling unified modeling of temporal flow-guided spatiotemporal dependencies. For the decoder part, a dynamic update module (DUM) is implemented to mitigate the “ghosting effect” by contextual information decay during upsampling. In the DUM, the initial temporal state is employed as gating units to reweight spatiotemporal features, refining multi-scale representations while preserving critical cloud motion signatures. Finally, we publicly develop and release a novel ground-based cloud image sequence dataset, comprising large-scale, high-resolution sequences captured under diverse meteorological conditions. The main contributions of this article are summarized as follows.
@@
    A Unified Spatiotemporal Module (USTM) is designed to serve as the core of the network. It features a Spatial Selection Module (SSM) for dynamic, adaptive multi-scale feature extraction and a low-complexity Temporal Agent Attention Module (TAM) that effectively balances predictive accuracy with computational efficiency.
@@
    A Dynamic Update Module (DUM) is introduced in the decoder. It leverages initial temporal states as an attention operator to reweight upsampled features, preserving critical motion signatures and effectively mitigating the “ghosting effect.”
@@
    The introduction and public release of the ASI-Cloud Image Sequence (ASI-CIS) dataset. As a major contribution to the community, ASI-CIS is a newly introduced, large-scale, high-resolution, multi-seasonal benchmark that addresses key limitations of previous datasets. It offers a more realistic and challenging foundation for advancing ground-based cloud extrapolation models. Extensive experiments on ASI-CIS show that USF-Net outperforms state-of-the-art methods in both prediction accuracy and computational efficiency.
The rest of this paper is organized as follows. In Section 2, we introduce the related works of ground-based cloud image sequence precidtion. In Section 3, we introduce the detailed structure of the proposed method. In Section 4, we evaluate the performance of our proposed methods. Finally, Section 5 concludes the paper.
@@ 2 Related Works
Accurate prediction of cloud image sequence extrapolation plays a pivotal role in enhancing the operational stability of grid-connected PV systems. Current methodologies are broadly divided into two categories: traditional methods and deep learning-based methods.
@@ 2.1 Traditional methods for cloud image sequence extrapolation
Conventional approaches predominantly employ optical flow (OF) algorithms for cloud image sequence extrapolation. OF is utilized to estimate the instantaneous velocity field of pixel-wise motion in ground-based cloud images, capturing spatiotemporal trends of cloud dynamics. Several studies have adopted OF-based methods for ground-based remote sensing cloud image sequence extrapolation. Omnidirectional optical flow tracking frameworks are proposed to establish quantitative relationships between cloud motion directionality and velocity magnitudes in cloud dynamics [Re1, Re2]. Boundary information within cloud imagery constitutes a critical determinant of prediction accuracy. Chang et al. implemented the Horn-Schunck OF algorithm to compute velocity variations for each pixel, augmenting motion field estimation through supplementary information integration [Re3]. However, these methods incur substantial computational overhead. Conversely, Wang et al. introduced a mathematical analysis framework to characterize inter-frame disparities, achieving reduced computational resource consumption while maintaining predictive performance [Re4].
@@ 2.2 DL-based methods for cloud image sequence extrapolation
In the historical development of DL, recurrent neural networks and their variant long short-term memory networks have served as pioneering methods for temporal sequence prediction. Several studies have successfully employed LSTM architecture to capture long-range dependencies in spatiotemporal sequences. However, standalone LSTM models exhibit elevated computational costs while struggling to extract complex spatial information effectively. Consequently, numerous works integrate CNN with LSTM frameworks to jointly extract localized spatial features and temporal information for sequential motion prediction. PredRNN enhanced the ConvLSTM architecture by introducing a zigzag memory flow mechanism to model short-term spatiotemporal dynamics [PredRNN]. Subsequently, PredRNN++ incorporated a gradient highway unit (GHU) to mitigate the gradient vanishing issue prevalent in LSTM-based models while integrating Causal-LSTM modules to strengthen spatial feature representation and short-term temporal modeling [PredRNN++]. Li et al. proposed cascaded Causal-LSTM layers to improve short-term prediction accuracy for cloud imagery. The model was augmented by GHUs with auxiliary skip connections to enhance spatiotemporal uniformity in modeling [CloudPredRNN++]. Nevertheless, ground-based remote sensing cloud image exhibits high-resolution cloud formations with multi-scale variations under complex meteorological conditions. Existing methodologies remain suboptimal for cloud sequence extrapolation tasks. To capture the dynamic states of cloud clusters at varying scales within cloud imagery, several studies have integrated multi-scale convolutional kernels. For instance, MSSTNet employs 3D convolutions with diverse kernel sizes to enhance the capacity of the model for multi-resolution image forecasting [Re5]. Wang et al. introduced 3D tensor augmentations within LSTM architectures to expand the effective receptive field [Re6]. However, the adoption of 3D convolution operations incurs significant computational overhead. The MSTANet employs multi-scale large kernels to aggregate multi-scale contextual information from cloud imagery while leveraging depthwise separable convolutions to construct large kernels with reduced computational complexity [MSTANet]. Accurate cloud image sequence extrapolation serves as a critical factor in ultra-short-term PV power forecasting, and the modeling of long-range spatiotemporal dependencies becomes particularly paramount. To this end, Chang et al. designed an MAU that simultaneously enlarges the model’s receptive field and captures spatial motion patterns across cloud sequences [MAU]. The Motion RNN introduces a Motion RGU module to unify transient variation modeling and motion trend representation [Re7]. When embedded within RNN architectures, this approach significantly improves spatiotemporal prediction accuracy under complex meteorological scenarios.
@@
Attention mechanisms have been demonstrated as an effective approach for establishing long-range dependencies, facilitating the extraction of temporal features in sequence prediction tasks. The STANet introduces a context gating unit (CGU) as an attention mechanism to unify the modeling of instantaneous cloud characteristics and motion trends [STANet]. Similarly, SimvPV2 incorporates a gated spatiotemporal attention module to enforce spatiotemporal consistency in sequence modeling [Re8]. Tan et al. decomposed temporal attention into intra-frame static attention and inter-frame dynamic attention through a dedicated temporal attention unit, capturing spatial features and temporal correlations [TAU]. Li et al. integrated a self-attention memory unit into the Cascaded Causal LSTM (CCLSTM) framework to extract long-term dependencies, enabling spatiotemporal modeling for cloud sequence prediction [CCLSTM]. Furthermore, the MSTANet introduces a multi-scale temporal attention mechanism that combines local temporal variations and global temporal variations, significantly enhancing the network’s temporal modeling capacity for cloud image extrapolation tasks [MSTANet].
@@
Unfortunately, while the methods above demonstrate commendable performance in cloud image sequence extrapolation tasks, they encounter three notable limitations. First, these approaches solely employ multi-scale convolutional kernels to capture contextual information, lacking adaptive mechanisms to extract features at varying resolutions dynamically. In addition, during spatiotemporal dependency modeling, the absence of temporal guidance hinders the unified integration of spatial and temporal information, resulting in suboptimal long-term dependency capture. Furthermore, existing attention mechanisms for temporal feature extraction neglect to balance computational complexity with prediction accuracy, leading to inefficiencies in practical deployment.
3 Proposed Method
This section provides a detailed exposition of the USF-Net architecture, beginning with the problem formulation and a high-level overview, followed by in-depth descriptions of its novel components: the Unified Spatiotemporal Module (USTM), the Dynamic Update Module (DUM), and the composite loss function.
@@ 3 Proposed Method
2
31ex
4repeat
5   Let layer $i$ = 1, lossloss = 0.0
6 $L=\lambda_{1}L_{M}+\lambda_{2}L_{MS}+\lambda_{3}L_{C}$ ($L$: loss function)
7 for $i\leftarrow 1$ to 33 do
@@ 6 $L=\lambda_{1}L_{M}+\lambda_{2}L_{MS}+\lambda_{3}L_{C}$ ($L$: loss function)
8    if $i=3$ then
9       $X_{T}\leftarrow X_{B}$
10      end if
11
12   end for
13
14 $X_{S}\leftarrow\text{SiB}(X_{B})$
15 $X_{T}\leftarrow\text{TiB}(X_{T})$
16 $X_{D}\leftarrow\text{DSM}(X_{S},X_{T})$
17 $D_{4}\leftarrow\text{DUM}(X_{D},X_{T_{0}})$
18 for $k\leftarrow 3$ to 11 do
@@ 17 $D_{4}\leftarrow\text{DUM}(X_{D},X_{T_{0}})$
19
20   end for
21 $Y\leftarrow\text{Conv}_{1\times 1}(D_{1})$
22 $\mathcal{L}\leftarrow\lambda_{1}L_{M}+\lambda_{2}L_{MS}+\lambda_{3}L_{C}$

### v2

@@ Abstract
Accurate extrapolation of ground-based cloud imagery is foundational to the stability of photovoltaic (PV) power generation systems. However, current methods rely on static kernels that neglect multi-scale cloud dynamics and lack robust temporal guidance for long-range dependency modeling. Moreover, the quadratic computational cost of standard attention mechanisms impedes real-time deployment. To bridge these gaps, we propose the Unified Spatiotemporal Fusion Network (USF-Net), which integrates adaptive large-kernel convolutions with a low-complexity attention mechanism. Specifically, USF-Net incorporates a unified spatiotemporal module (USTM), which comprises a spatial information branch equipped with a spatial selection module (SSM) for dynamic multi-scale context extraction and a temporal information branch featuring a temporal agent attention module (TAM). The TAM efficiently models long-range temporal dependencies, circumventing the computational bottlenecks of traditional attention. In the decoder, a dynamic update module (DUM) leverages initial temporal states to preserve motion signatures, thereby mitigating ghosting effects. Crucially, we introduce the ASI-Cloud Image Sequence (ASI-CIS) dataset, a large-scale, high-resolution benchmark designed to address current data limitations. Extensive experiments on ASI-CIS demonstrate that USF-Net establishes a new state-of-the-art, offering a superior trade-off between prediction accuracy and computational efficiency for ground-based cloud forecasting. The dataset and source code will be available at https://github.com/she1110/ASI-CIS.
@@
Unified spatiotemporal, Ground-based cloud images, Temporal guidance, Spatiotemporal prediction, Photovoltaic power prediction
@@ 1 Introduction
Driven by the imperative to decarbonize the global energy landscape, photovoltaic (PV) power generation has established itself as a cornerstone of renewable energy transitions, distinguished by its zero-carbon emissions and ubiquitous resource availability [Intro1]. However, the integration of PV systems into the electrical grid is complicated by the stochastic nature of solar irradiance. Rapid fluctuations in power output, frequently precipitated by highly dynamic cloud cover, impose severe stress on power dispatch systems and necessitate robust energy storage configurations to maintain grid stability [Intro2]. Consequently, high-resolution solar irradiance forecasting has become a prerequisite for enhancing PV integration capacity and ensuring power system reliability.
@@
Since PV output exhibits a strong correlation with solar irradiance, which is primarily modulated by atmospheric cloud dynamics, precise cloud observation is critical [Intro3]. Irradiance variability is generally categorized into sustained shading (e.g., extensive stratiform coverage) and transient shading (e.g., rapid cumulonimbus convection). While numerical weather prediction (NWP) models effectively forecast sustained events, they lack the temporal granularity required for transient shading [Intro4]. Conversely, satellite imagery, though offering broad coverage, typically suffers from coarse spatial resolution (>>1 km) and high latency ($\geq$30 minutes), rendering it inadequate for tracking localized, micro-scale cloud evolution [Intro5, Intro6]. In contrast, ground-based sky imagers capture cloud structure and motion with high spatiotemporal resolution, making them the optimal modality for ultra-short-term irradiance forecasting [25TIM]. Therefore, the accurate extrapolation of ground-based cloud image sequences, essentially a deterministic spatiotemporal prediction problem, is pivotal for operational grid stability [Intro7, Intro8, Intro9].
@@
Advancing the state-of-the-art (SOTA) in cloud extrapolation requires addressing two fundamental challenges: (a) modeling the complex, nonlinear, and multi-scale morphological deformations of clouds, and (b) efficiently capturing long-range spatiotemporal dependencies to satisfy the real-time inference constraints of industrial applications.
@@
Existing methodologies fall into two primary paradigms: traditional optical flow frameworks and deep learning-based approaches. Traditional methods, often relying on vector field representations and similarity maximization, are frequently hampered by excessive computational overhead and limited capacity to represent nonlinear features, leading to performance degradation under complex atmospheric conditions [Intro10, Intro11].
The advent of deep learning has shifted the paradigm toward data-driven spatiotemporal modeling. Recurrent Neural Networks (RNNs)  [Intro12] and their variants, such as Long Short-Term Memory (LSTM) networks [Intro14], have been widely adopted for their ability to model temporal dynamics. To address the spatial limitations of standard RNNs, hybrid architectures like the Convolutional LSTM (ConvLSTM) were introduced to integrate spatial feature extraction with temporal memory [ConvLSTM]. Subsequent innovations have focused on multi-scale convolutional kernels to capture hierarchical cloud morphology and attention mechanisms to model global dependencies [Intro16, Intro17].
Despite these advancements, significant limitations persist. First, the reliance on static, fixed-size kernels prevent models from dynamically adapting their receptive fields (RF) to the diverse scales of cloud structures. Second, the interaction between spatial and temporal feature streams is often decoupled, lacking an explicit mechanism for temporal flow to guide spatial feature learning. Third, the quadratic complexity of standard self-attention mechanisms often creates a computational bottleneck prohibitive for high-resolution, real-time monitoring. Finally, the ghosting effect, a blurring artifact prevalent in prediction decoders, remains a critical issue, often exacerbated by the decay of contextual information during upsampling.
Furthermore, current research is constrained by the quality of available benchmarks. Existing datasets often suffer from low spatial resolution and hardware-induced visual artifacts, which introduce noise and limit the applicability of extrapolation models to real-world PV forecasting [MSTANet, CloudPredRNN++]. Thus, there is a pressing need for a high-resolution, multi-scale, and cross-seasonal benchmark dataset.
@@
To bridge these gaps, this article proposes the Unified Spatiotemporal Fusion Network (USF-Net), a novel encoder-decoder architecture that integrates adaptive large-kernel convolutions with a low-complexity attention mechanism. Specifically, the encoder employs depthwise separable (DW) convolutions and squeeze-excitation (SE) blocks for efficient multi-scale feature extraction. Following the encoder, we propose a unified spatiotemporal module (USTM) comprising three core components: 1) a spatial information branch, in which a spatial selection module (SSM) is designed for dynamic receptive field adjustment; 2) a temporal information branch incorporating a temporal agent attention module (TAM) is introduced for efficient dependency modeling; and 3) a dynamic spatiotemporal module: a temporal guidance module (TGM) is proposed to explicitly fuse temporal flow with spatial features. For the decoder part, a dynamic update module (DUM) is implemented to mitigate the ghosting effect by using the initial temporal state as a gating operator to reweight spatiotemporal features while preserving high-frequency motion signatures. Additionally, we introduce and publicly release the ASI-CIS dataset, a large-scale, high-resolution benchmark for ground-based cloud extrapolation. The main contributions of this work are summarized as follows:
@@
    A USTM is designed to serve as the core of the network. It features a SSM for dynamic, adaptive multi-scale feature extraction and a low-complexity TAM that effectively balances predictive accuracy with computational efficiency.
@@
    A DUM is introduced in the decoder. It leverages initial temporal states as an attention operator to suppress the ghosting effect and preserve fine-grained cloud details.
@@
    The introduction and public release of the ASI-CIS dataset. As a major contribution to the community, ASI-CIS is a newly introduced, large-scale, high-resolution, multi-seasonal benchmark that addresses key limitations of previous datasets. It offers a more realistic and challenging foundation for advancing ground-based cloud extrapolation models. Extensive experiments on ASI-CIS show that USF-Net outperforms state-of-the-art methods in both prediction accuracy and computational efficiency.
The remainder of this paper is organized as follows. In Section 2, we introduce the related works of ground-based cloud image sequence precidtion. In Section 3, we introduce the detailed structure of the proposed method. In Section 4, we evaluate the performance of our proposed methods. Finally, Section 5 concludes the paper.
@@ 2 Related works
Precise cloud image sequence extrapolation underpins the operational stability of grid-connected PV systems. Existing methodologies categorize primarily into two categories: traditional methods and deep learning-based methods.
@@ 2.1 Traditional methods for cloud image sequence extrapolation
Traditional frameworks predominantly leverage optical flow (OF) algorithms. These techniques estimate the instantaneous velocity field of pixel-wise motion in ground-based cloud images, capturing spatiotemporal trends of cloud dynamics. Several studies have adopted OF-based methods for ground-based remote sensing cloud image sequence extrapolation. Omnidirectional tracking frameworks quantify relationships between cloud motion directionality and velocity magnitudes in cloud dynamics [Re1, Re2]. Boundary information within cloud imagery constitutes a critical determinant of prediction accuracy. Chang et al. implemented the Horn-Schunck OF algorithm to compute velocity variations for each pixel, augmenting motion field estimation through supplementary information integration [Re3]. Nevertheless, such methods impose significant computational burdens. Conversely, Wang et al. introduced a mathematical analysis framework to characterize inter-frame disparities, achieving reduced computational resource consumption while maintaining predictive performance [Re4].
@@ 2.2 DL-based methods for cloud image sequence extrapolation
Historically, recurrent neural networks and their variant long short-term memory networks constituted the foundational approach for temporal sequence prediction. Several studies have successfully employed LSTM architecture to capture long-range dependencies in spatiotemporal sequences. However, standalone LSTM models exhibit elevated computational costs while struggling to extract complex spatial information effectively. Consequently, numerous works integrate CNN with LSTM frameworks to jointly extract localized spatial features and temporal information for sequential motion prediction. Subsequent architectures augmented the standard Convolutional LSTM by introducing a bidirectional memory propagation mechanism to model short-term spatiotemporal dynamics [PredRNN]. Subsequently, Further iterations incorporated gradient highway units to alleviate vanishing gradients in LSTM-based models while integrating Causal-LSTM modules to strengthen spatial feature representation and short-term temporal modeling [PredRNN++]. Li et al. proposed cascaded Causal-LSTM layers to improve short-term prediction accuracy for cloud imagery. The model was augmented by GHUs with auxiliary skip connections to enhance spatiotemporal uniformity in modeling [CloudPredRNN++]. However, ground-based remote sensing cloud imagery characterizes by high-resolution cloud formations with multi-scale variations under complex meteorological conditions. Existing methodologies remain suboptimal for cloud sequence extrapolation tasks. To resolve multi-scale dynamic states of cloud clusters at varying scales within cloud imagery, several studies have integrated multi-scale convolutional kernels. For instance, MSSTNet employs 3D convolutions with diverse kernel sizes to enhance the capacity of the model for multi-resolution image forecasting [Re5]. Wang et al. introduced 3D tensor augmentations within LSTM architectures to expand the effective receptive field [Re6]. However, the adoption of 3D convolution operations incurs significant computational overhead. The MSTANet employs multi-scale large kernels to aggregate multi-scale contextual information from cloud imagery while leveraging depthwise separable convolutions to construct large kernels with reduced computational complexity [MSTANet]. Accurate cloud image sequence extrapolation serves as a critical factor in ultra-short-term PV power forecasting, and the modeling of long-range spatiotemporal dependencies becomes particularly paramount. To this end, Chang et al. designed an MAU that simultaneously enlarges the model’s receptive field and captures spatial motion patterns across cloud sequences [MAU]. The Motion RNN introduces a Motion RGU module to unify transient variation modeling and motion trend representation [Re7]. When embedded within RNN architectures, this approach significantly improves spatiotemporal prediction accuracy under complex meteorological scenarios.
@@
Attention mechanisms have proven effective for establishing long-range dependencies, facilitating the extraction of temporal features in sequence prediction tasks. The STANet introduces a context gating unit (CGU) as an attention mechanism to unify the modeling of instantaneous cloud characteristics and motion trends [STANet]. Similarly, SimvPV2 incorporates a gated spatiotemporal attention module to enforce spatiotemporal consistency in sequence modeling [Re8]. Tan et al. decomposed temporal attention into intra-frame static attention and inter-frame dynamic attention through a dedicated temporal attention unit, capturing spatial features and temporal correlations [TAU]. Li et al. integrated a self-attention memory unit into the Cascaded Causal LSTM (CCLSTM) framework to extract long-term dependencies, enabling spatiotemporal modeling for cloud sequence prediction [CCLSTM]. Furthermore, the MSTANet introduces a multi-scale temporal attention mechanism that combines local temporal variations and global temporal variations, significantly enhancing the network’s temporal modeling capacity for cloud image extrapolation tasks [MSTANet].
@@
While demonstrating efficacy, these methodologies encounter three distinct limitations. First, these approaches solely employ multi-scale convolutional kernels to capture contextual information, lacking adaptive mechanisms to extract features at varying resolutions dynamically. In addition, during spatiotemporal dependency modeling, the absence of temporal guidance hinders the unified integration of spatial and temporal information, resulting in suboptimal long-term dependency capture. Furthermore, existing attention mechanisms for temporal feature extraction neglect to balance computational complexity with prediction accuracy, leading to inefficiencies in practical deployment.
@@
1ex
repeat
   Let layer $i$ = 1, lossloss = 0.0
 $L=\lambda_{1}L_{M}+\lambda_{2}L_{MS}+\lambda_{3}L_{C}$ ($L$: loss function)
 for $i\leftarrow 1$ to 33 do
@@
    if $i=3$ then
       $X_{T}\leftarrow X_{B}$
      end if
   end for
 $X_{S}\leftarrow\text{SiB}(X_{B})$
 $X_{T}\leftarrow\text{TiB}(X_{T})$
 $X_{D}\leftarrow\text{DSM}(X_{S},X_{T})$
 $D_{4}\leftarrow\text{DUM}(X_{D},X_{T_{0}})$
 for $k\leftarrow 3$ to 11 do
@@
   end for
 $Y\leftarrow\text{Conv}_{1\times 1}(D_{1})$
 $\mathcal{L}\leftarrow\lambda_{1}L_{M}+\lambda_{2}L_{MS}+\lambda_{3}L_{C}$
@@
3 Proposed Method
This section elucidates the architecture of USF-Net, commencing with the mathematical formulation of the extrapolation task. Subsequently, we detail the novel architectural components: the unified spatiotemporal module, the dynamic update module, and the composite loss function designed to enforce structural fidelity.